Hi, thanks so so much for helping me out!!! I am working on substitution and am really confused! here is my problem
-5x+3y=12
x+2y=8
I Tried:
-5x+3y=12
x+2y=8
(-5)+2y=8
+5 +5 But then I was told I need 2 answers and I don't know how
2y= 13 to find the second answer!
13/2
y =6.5


Found 2 solutions by rfer, cleomenius:
Answer by rfer(16322)

x+2y=8
change to
x=-2y+8
------------
-5x+3y=12
-5(-2y+8)+3y=12
10y-40+3y=12
13y=52
y=4
------------------
-5x+3*4=12
-5x+12=12
-5x=12-12
x=0
---------------
x+2y=8
0+2*4=8
-----------------
Does that help?
Bob

Answer by cleomenius(959)

I set the second equation up as x = -2y + 8
Then substituted into the first equation:
-5(-2y + 8) + 3y = 12
10y -40 + 3y = 12
13y = 52
y = 4
into first equation:
-5x + 12 = 12 , x = 0
into second equation
0 + 2(4) = 8.
Cleomenius.